2 triangles are similar
Triangle RTS where rt=4.2 and st=6.0 do not know rs
Triangle JLK where lk=8
what is the length of JL

Answer by Alan3354(69443)

Since they're similar, the ratios are the same. 6:8::4.2:JL
Multiply means and extremes - 6*JL = 8*4.2
6*JL = 33.6
JL = 5.6